The escape velocity of projection from the earth is approximately (R = 6400 km)
(a) 7 km/sec
(b) 112 km/sec
(c) 12.2 km/sec
(d) 1.1 km/sec

Grade:12

1 Answers

Dakshita Sharma
5 years ago
1/2 mv^2=GM/Rv=Since,1/2mv^2 = GM/Rv=√(2GM/R)Since,g=GM/R^2v=√(2gR)v=√(2×9.8×6.4×10^6v=11.2×10^3 m/sv=11.2 km/s